# IterNorm: Fast Iterative Normalization

### v1

@@ V-D LLM-level evaluation
TABLE IV: LLM-level evaluation of IterNorm using OPT-125M on two text generation tasks
Format Baseline perplexity IterNorm # steps Perplexity # steps Perplexity WikiText-2 FP32 FP16 BFloat16 18.21 25.35 19.17 3 18.37 (+0.16) 25.51 (+0.16) 19.43 (+0.26) 5 18.21 (+0.00) 25.35 (+0.00) 19.20 (+0.03) FP32 FP16 BFloat16 18.21 25.35 19.17 4 18.22 (+0.01) 25.35 (+0.00) 19.20 (+0.03) 10 18.21 (+0.00) 25.35 (+0.00) 19.17 (+0.00) BST FP32 FP16 BFloat16 17.30 19.61 17.83 3 17.36 (+0.06) 19.67 (+0.06) 17.91 (+0.08) 5 17.30 (+0.00) 19.61 (+0.00) 17.84 (+0.01) FP32 FP16 BFloat16 17.30 19.61 17.83 4 17.31 (+0.01) 19.61 (+0.00) 17.84 (+0.01) 10 17.30 (+0.00) 19.61 (+0.00) 17.83 (+0.00)
We evaluated IterNorm on the LLM-level by creating a PyTorch module for IterNorm. Note that the precision of IterNorm in PyTorch negligibly differs from that in FPGA. We considered Open Pre-trained Transformer model with 125M parameters (OPT-125M) [14] for text generation. The OPT-125M model is a decoder-only model that consists of a stack of 12 transformer blocks, each of which employs 12 attention heads with an embedding size of 768. We considered two text generation datasets: WikiText-2 [15] and Blended Skill Talk (BST) [16]. We replaced all layer normalization blocks in the OPT-125M pre-trained on each dataset by IterNorm and measured a change in the perplexity score for text generation as the LLM-level error of IterNorm. The IterNorm module takes the number of iteration steps $n_{\text{iter}}$ as a parameter. The perplexity scores for different iteration steps in FP32, FP16, and BFloat16 are listed in Table IV. Compared to the baseline perplexity, the perplexity scores for both WikiText-2 and BST marginally increased after the third iteration step.
@@ VI Related work and comparison
Our IterNom macros designed for FP32, FP16, and BFloat16 are compared with previous implementations of layer normalization in Table III. The method in [8] realizes the layer normalization of INT32 vectors using integer-only arithmetic. To this end, they adopted an iterative algorithm [17] that approximates square root values. Thus, this method requires additional division operations to normalize the input integer vector. When implemented in 65nm CMOS, the area and power overheads of the circuit are 68.3 mm2 and 2.0 W, respectively.
@@ VII Conclusion
We introduced IterNorm, an efficient method for iteratively L2-normalizing input vectors without costly division or square root operations. Grounded in solid theory, IterNorm is applicable to general FP data and ensures high precision, outperforming FISR in six out of nine cases for FP32 and five out of nine for BFloat16 across the embedding lengths used in the OPT models. It also converges quickly, reaching its fixed point within five iterations. Implemented in 32/28nm CMOS technology, the IterNorm macro processes $d$-dimensional input vectors, where $64\leq d\leq 1024$, with a latency of 112-227 clock cycles.

### v2

@@ V-D LLM-level evaluation
TABLE IV: LLM-level evaluation of IterL2Norm using OPT-125M and OPT-350M on two text generation tasks
Task Format OPT-125M OPT-350M Baseline # steps Perplexity Baseline # steps Perplexity Wikitext-2 FP32 18.21 3 18.37 (+0.16) 15.28 3 15.28 (+0.00) 4 18.22 (+0.01) 4 15.28 (+0.00) 5 18.21 (+0.00) 5 15.28 (+0.00) 10 18.21 (+0.00) 10 15.28 (+0.00) FP16 25.35 3 25.51 (+0.16) 27.57 3 27.57 (+0.00) 4 25.35 (+0.00) 4 27.57 (+0.00) 5 25.35 (+0.00) 5 27.57 (+0.00) 10 25.35 (+0.00) 10 27.57 (+0.00) BFloat16 19.17 3 19.43 (+0.26) 15.43 3 15.44 (+0.01) 4 19.20 (+0.03) 4 15.43 (+0.00) 5 19.20 (+0.03) 5 15.43 (+0.00) 10 19.17 (+0.00) 10 15.43 (+0.00) BST FP32 17.30 3 17.36 (+0.06) 15.41 3 15.41 (+0.00) 4 17.31 (+0.01) 4 15.41 (+0.00) 5 17.30 (+0.00) 5 15.41 (+0.00) 10 17.30 (+0.00) 10 15.41 (+0.00) FP16 19.61 3 19.67 (+0.16) 21.94 3 21.95 (+0.01) 4 19.61 (+0.00) 4 21.94 (+0.00) 5 19.61 (+0.00) 5 21.94 (+0.00) 10 19.61 (+0.00) 10 21.94 (+0.00) BFloat16 17.83 3 17.91 (+0.08) 15.49 3 15.49 (+0.00) 4 17.84 (+0.01) 4 15.49 (+0.00) 5 17.84 (+0.01) 5 15.49 (+0.00) 10 17.83 (+0.00) 10 15.49 (+0.00)
We evaluated IterL2Norm at the LLM-level by creating a PyTorch module for IterL2Norm. Note that the precision of IterL2Norm in PyTorch negligibly differs from that in FPGA. We considered the Open Pre-trained Transformer (OPT) models with 125M (OPT-125M) and 350M (OPT-350M) parameters [14] for text generation. The OPT models are decoder-only models consisting of stacks of 12 and 24 transformer blocks, each employing 12 and 16 attention heads with embedding sizes of 768 and 1024, respectively. We used two text generation datasets: WikiText-2 [15] and Blended Skill Talk (BST) [16]. All layer normalization blocks in the pre-trained OPT-125M and OPT-350M models for each dataset were replaced with IterL2Norm, and we measured the change in perplexity scores for text generation as the LLM-level error of IterL2Norm. The IterL2Norm module takes the number of iteration steps $n_{\text{iter}}$ as a parameter. The perplexity scores for different iteration steps in FP32, FP16, and BFloat16 are listed in Table IV. Compared to the baseline perplexity, the perplexity scores for both WikiText-2 and BST marginally increased after the third iteration step.
@@ VI Related work and comparison
Our IterL2Nom macros designed for FP32, FP16, and BFloat16 are compared with previous implementations of layer normalization in Table III. The method in [8] realizes the layer normalization of INT32 vectors using integer-only arithmetic. To this end, they adopted an iterative algorithm [17] that approximates square root values. Thus, this method requires additional division operations to normalize the input integer vector. When implemented in 65nm CMOS, the area and power overheads of the circuit are 68.3 mm2 and 2.0 W, respectively.
@@ VII Conclusion
We introduced IterL2Norm, an efficient method for iteratively L2-normalizing input vectors without costly division or square root operations. Grounded in solid theory, IterL2Norm is applicable to general FP data and ensures high precision, outperforming FISR in six out of nine cases for FP32 and five out of nine for BFloat16 across the embedding lengths used in the OPT models. It also converges quickly, reaching its fixed point within five iterations. Implemented in 32/28nm CMOS technology, the IterL2Norm macro processes $d$-dimensional input vectors, where $64\leq d\leq 1024$, with a latency of 116-227 clock cycles.